NUCLEAR SPIN OF H_{3}^{+} AND H_{2} IN DENSE MOLECULAR CLOUDS

\underline{KYLE N. CRABTREE}, Department of Chemistry, University of Illinois, Urbana, IL 61801; BEN-JAMIN J. McCALL, Departments of Chemistry, Astronomy, and Physics, University of Illinois, Urbana, IL 61801.

The ortho:para ratio of H_{2} is a critical parameter for deuterium fractionation in cold, dense quiescent cores. The dominant reservoir for interstellar deuterium is in the inert molecule HD, but the exothermic reaction H_{3}^{+} + HD → H_{2}D^{+} + H_{2} + 220 K (and H_{2}D^{+} + HD, etc.) can yield highly reactive species capable of distributing deuterium to other molecules. The barrier to the reverse reaction, however, can be overcome even at temperatures below 10 K when ortho-H_{2} (o-H_{2}) reacts with H_{2}D^{+} (or D_{2}H^{+}, D_{3}^{+}), as ortho-H_{2} possesses ∼170 K of internal rotational energy in its ground state. Recent modeling work has demonstrated the importance of o-H_{2} in cold, dense, highly depleted cores using a chemical network that includes all nuclear spin modifications of H_{3}^{+}, H_{2}, and their isotopologues, but the initial o-H_{2} fraction is taken as a parameter in the model. Observationally or computationally constraining this quantity would aid in understanding deuterium fractionation in dense cores. To learn about the initial o-H_{2} fraction in a cold core, we have modeled the chemistry of non-depleted dense interstellar clouds from which cold cores are thought to form. A simplified gas-phase chemical network consisting of 28 species and ∼170 reactions is combined with a physical model of a dense cloud, including time-dependent physical conditions. Included in the network are the nuclear spin modifications of H_{2}, H_{2}^{+}, and H_{3}^{+}, as well as nuclear spin dependent rate coefficients for the thermalization reactions H_{2} + H^{+} and H_{3}^{+} + H_{2}. By modeling the time-dependent chemistry, we find that the ortho:para ratio of H_{2} requires 10^{7}-10^{8} years to reach steady state under “standard” dense cloud conditions, which is at least on the order of the cloud lifetime. The timescale depends on the ionization rate, the rate coefficients of the various H_{3}^{+} + H_{2} reactions, and the relative abundances of H_{3}^{+} and H^{+}, but is largely insensitive to the total density and temperature. Even at steady state, the o-H_{2} fraction is calculated to be >0.5% at 10 K, which is several orders of magnitude above its value at thermodynamic equilibrium. The prospects for using observations of the ortho:para ratio of H_{3}^{+} as a probe of the H_{2} ortho:para ratio will be discussed.